6. Calculate the pH of a solution that is 0.60 Min HF (Ka = 7.2 x 10 ) and 0.80 Min NaF.
Answer #1

Concentration of HF = 0.60M

Concentration of NaF = 0.80M

Ka = 7.2x10^-4

-log(Ka) = -log(7.2x10^-4)

PKa = 3.14

PH = Pka + log[salt]/[acid]

PH = 3.14 +log(0.80/0.60)

PH = 3.26

PH of the solution = 3.26